[Brahms-l] Shift report 20040125 08:00-16:00

From: Apache <apache@bnl.gov>
Date: Sun Jan 25 2004 - 16:05:29 EST
------------------------------------------------------------
Date: Sun Jan 25 16:05:29 2004, Shift: 08:00-16:00
Supervisor: fv, with: 
Log book#15 pages 42 - 42

Shift Summary:
------------------------------------------------------------
Shift summary

-- beam from 10:00 until 15:30 where dump was requested by 3/4. -> 5.5 hours

-- BFS platfomr cannot be moved. Taked to Dana. We will request access tomorrow for repair.

-- until then take data for
10 deg (B306) 1M and swith POL A306 (1M). Thus should last until tomorrow.

-- Spend part of shift doing trigger studies; a few dedicated runs with single triggers.
-- Trigger timing for FS was adjusted after 9857.
ZDC width set to 4 nsec rather than 10
- implemented trigger 5 as prev setup by Dana
i.e. ZDC vertex trigger ~ effecient from +-22 cm > 90%.

special runs:
9857 trigger 6 only  81K
9858  trigger 6 only 54K
9859 trigger 2 only  58K
9860 trigger 6 only  59K (lower HV 1650 on trFS)
9861 trigger 4 only 124K (min bias to checkout effeciencies

10 deg B306; 40 deg A2255
9862 120K tr6; 114K tr 3 
9863 75K tr6 ; 70K tr 3
9864 53K tr 6; 41K  tr 3
9865 69k tr 6; 53 tr 3; Add trigger 5
9866 72K tr 6; 56k tr4;

B306 == 370K so far; 
--problems
BFS cannot move
The beam monitor does not update fill no and fille time, while current+mom is ok.

see also trigger logbook page 92/93
------------------------------------------------------------

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Sun Jan 25 16:06:18 2004

This archive was generated by hypermail 2.1.8 : Sun Jan 25 2004 - 16:06:33 EST